OpenStudy (anonymous):

sove by x^2-6x+3=0 by completing the square?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

I think you mean "Solve for x^2-6x+3=0 by completing the square." So what you want to do is, you have that -6, and half of that is -3, so you write (x-3)*(x-3) (your job is to expand this now and see what you get)

OpenStudy (anonymous):

x=9 and x=3

OpenStudy (anonymous):

No I want you to expand (x-3)*(x-3). It should start with x^2...

OpenStudy (anonymous):

x^2-6x-9=-3-9

OpenStudy (anonymous):

close, I think you are trying to do too many steps at once but you have the right Idea. We know that (x-3)*(x-3) = x^2 - 6x + 9. We want to solve for x^2 - 6x + 3 = 0 so now we do the old plus and minus trick... = x^2 - 6x + 3 = 0 = x^2 - 6x + 9 - 9 + 3 = 0 and now we can substitute in the equation on top... = (x-3)*(x-3) - 9 + 3 = 0 = (x-3)*(x-3) - 6 = 0 (x-3)*(x-3) = 6 = (x-3)^2 = 6 (x-3) = +-sqrt(6) x = 3 +- sqrt(6) where "+-" means "plus or minus"

OpenStudy (anonymous):

so is x=9 and x=3 right or no i didnt quite understand what u meant